Please help:
The length of a rectangular swimming pool is twice its width. The pool is surrounded by a walk that is 2 feet wide. The area of the region consisting of the pool and the walk is 1056 square feet.
(a) Use the method of completing the square to determine the dimensions of the swimming pool.
The length of the pool is ____ feet, and the width of the pool is ____feet.
(b) If the material for the walk costs $8 per square foot, how much would the material cost for the entire walk?
The material would cost $____ .

Answer by jorel555(1290)

Say the width of your pool is w; the length is then equivalent to 2w. The width and length of your walkway are w+4 (2 feet are added to each side, thus adding 4 to the total), and 2w+4. Thus the area is:
(w+4)(2w+4)=1056
2w^2+12w+16=1056
2w^2+12w-1040=0
w^2+6w-520=0
(w+26)(w-20)=0
w=-26,20
Our only real solution is width=20, length=40!!!!!!!!